### v6

@@ 7 Conclusion
Finally, from a theoretical standpoint, we have discussed a type of communication that – we believe – is somewhat overlooked by the economic literature: the sort of persuasion that comes from the tangible results of economic actions, rather than from information design by a sender or the implicit informational content of a given equilibrium behavior. We believe that a deeper investigation of this force could help explain a variety of phenomena that go beyond productive incentives, including voter polarization, excessive debt accumulation, and other policy distortions.
@@ A.1 Examples
The characteristics of the views of example 3 – in particular, the supermodularity of expected output and noise log-concavity – recall the returns structure of Heidhues et al., (2018).
@@ A.3 Benchmark: Unawareness of Disagreement and Myopia
Next, consider part (ii). Let players be unaware of disagreement. Fix $i\in\{A,B\}$ and $m_{1}^{i}\in M$. By unawareness of disagreement, player $i$ plays as if the initial models of the two players were $(m^{i}_{1},m^{i}_{1})$. Since the two hypotheses compared by player $i$ at the end of $t=1$ coincide – the alternative model is the same as the null model – it has $m^{i}_{1}=m^{i}_{2}$ with certainty. Hence, if $m^{i}_{1}=\mathcal{H}$ then, $i$’s behavior is pinned down by 6. If $m^{i}_{1}=\mathcal{L}$ then, $i$’s behavior is pinned down by 7. It follows that
@@ A.4 Other Proofs
Consider now the case where there exists $k\in\{k^{A},k^{B}\}$ such that $m^{A}_{k}\neq m^{B}_{k}$. Without loss of generality, assume $k^{A}$ is such that $m^{A}_{k^{A}}\neq m^{B}_{k^{A}}$. Fix effort $e^{A},e^{B}\in\mathcal{E}$. $\Pi_{(e^{A},k^{A})}=(m^{A}_{k^{A}}(\cdot|e^{A}),m^{B}_{k^{A}}(\cdot|e^{A}),\mathcal{Y})$ is an informative experiment for states $m^{A}$ and $m^{B}$. There are two possible cases. First, if $m^{A}_{k^{B}}=m^{B}_{k^{B}}$, then $(y^{B},e^{B})$ are uninformative when it comes to discriminating between $m^{A}$ and $m^{B}$. Because $y^{A}$ and $y^{B}$ are independent, test LR will only use information about $(y^{A},e^{A},k^{A})$. Hence, we have $\phi_{m^{A}}^{m^{B}}(e^{A},e^{B},k^{A},k^{B}|m^{A})=\phi_{m^{A}_{k^{A}}}^{m^{B}_{k^{A}}}(e^{A},k^{A}|m^{A}_{k^{A}})$ which is increasing in $e^{A}$ by lemma A.1. The second case is $m^{A}_{k^{B}}\neq m^{B}_{k^{B}}$, in which case $\Pi_{(e^{B},k^{B})}=(m^{A}_{k^{B}}(\cdot|e^{B}),m^{B}_{k^{B}}(\cdot|e^{B}),\mathcal{Y})$ is also an informative experiment with states $m^{A}$ and $m^{B}$. Assume that this is the case and define the composite experiment $\Pi_{(e^{A},k^{A},e^{B},k^{B})}=(\Pi_{(e^{A},k^{A})},\Pi_{(e^{B},k^{B})})$. Next, consider the composite experiment $\Pi_{(\hat{e}^{A},k^{A},e^{B},k^{B})}=(\Pi_{(\hat{e}^{A},k^{A})},\Pi_{(e^{B},k^{B})})$, where $\hat{e}^{A}>e^{A}$. By assumption 3, $\Pi_{(\hat{e}^{A},k^{A})}$ is more informative than $\Pi_{(e^{A},k^{A})}$. Since $\Pi_{(\hat{e}^{A},k^{A})}$, $\Pi_{(e^{A},k^{A})}$ and $\Pi_{(e^{B},k^{B})}$ are independent, $\Pi_{(\hat{e}^{A},k^{A},e^{B},k^{B})}$ is more informative than $\Pi_{(e^{A},k^{A},e^{B},k^{B})}$ – the proof is analogous to the one provided for lemma 4.1. Hence, by lemma A.1, $\phi_{m^{A}}^{m^{B}}(\hat{e}^{A},e^{B},k^{A},k^{B}|m^{A})\geq\phi_{m^{A}}^{m^{B}}(e^{A},e^{B},k^{A},k^{B}|m^{A})$, which proves that $\phi_{m^{A}}^{m^{B}}(e^{A},e^{B},k^{A},k^{B}|m^{A})$ is increasing in $e^{A}$. The proof that $\phi_{m^{A}}^{m^{B}}(e^{A},e^{B},k^{A},k^{B}|m^{A})$ is increasing in $e^{B}$ is a replication of the same argument. The proof for $\phi_{m^{B}}^{m^{A}}(e^{A},e^{B},k^{A},k^{B}|m^{B})$ is increasing in $e^{A}$ and $e^{B}$ is analogous.
@@
Next, note that, if Ann and Bob’s initial models are $(\mathcal{H}_{x},\mathcal{L}_{y})$ and $(\mathcal{L}_{x},\mathcal{H}_{y})$, at $t=2$ both players will certainly have one view of type $\mathcal{H}$ – although they might still hold such view for different technologies. This means that, at $t=2$, both players will choose effort $e^{\mathcal{H}}$. Hence, by strict FOSD monotonicity and $Q_{x}=Q_{y}$,
@@
Next, we move to the proof of part (ii). Let $m\in M$. By the standard arguments of the previous proposition, when players start the game agreeing on $m$, they will exert at most effort $e^{\mathcal{H}}$ in every period. Since models $(\mathcal{H},\mathcal{L})$ and $(\mathcal{L},\mathcal{H})$ are both correct with probability $\frac{1}{2}$, and effort and technology choices are only contingent on models – which are not changed between periods – it must hold
$\mathbb{E}_{p}[\hat{Y}(m,m,Q)]\leq 2(y^{\mathcal{H}}_{\mathcal{H}}+y^{\mathcal{H}}_{\mathcal{L}}).$
Now, consider the disagreeing team with $\mathbf{m}_{1}=((\mathcal{H}_{x},\mathcal{L}_{y}),(\mathcal{L}_{x},\mathcal{H}_{y}))$. By our previous results, $s^{i}_{1}((\mathcal{H}_{x},\mathcal{L}_{y}),(\mathcal{L}_{x},\mathcal{H}_{y}))=e^{\mathcal{H}}+\hat{\Delta}_{e}$, with $\hat{\Delta}_{e}\geq 0$.
@@
$\hat{Y}(m^{A}_{1},m^{B}_{1},Q)=\mathbb{E}_{\mathcal{H}}[Y|e^{\mathcal{H}}+\hat{\Delta}_{e},x]+\mathbb{E}_{\mathcal{L}}[Y|e^{\mathcal{H}}+\hat{\Delta}_{e},y]+y^{\mathcal{H}}_{\mathcal{H}}$
@@
$\hat{Y}(m^{A}_{1},m^{B}_{1},Q)=\mathbb{E}_{\mathcal{H}}[Y|e^{\mathcal{H}}+\hat{\Delta}_{e},y]+\mathbb{E}_{\mathcal{L}}[Y|e^{\mathcal{H}}+\hat{\Delta}_{e},x]+y^{\mathcal{H}}_{\mathcal{H}}$
@@
$\mathbb{E}_{p}[\hat{Y}(m^{A}_{1},m^{B}_{1},Q)]=y^{\mathcal{H}}_{\mathcal{H}}+y^{\mathcal{H}}_{\mathcal{L}}+\mathbb{E}_{\mathcal{H}}[Y|e^{\mathcal{H}}+\hat{\Delta}_{e},x]+\mathbb{E}_{\mathcal{L}}[Y|e^{\mathcal{H}}+\hat{\Delta}_{e},y]$
@@
$\mathbb{E}_{p}[\hat{Y}(m^{A}_{1},m^{B}_{1},Q)]-\mathbb{E}_{p}[\hat{Y}(m,m,Q)]\geq\underbrace{(\mathbb{E}_{\mathcal{H}}[Y|e^{\mathcal{H}}+\hat{\Delta}_{e},x]-y^{\mathcal{H}}_{\mathcal{H}})}_{\geq 0}$

### v7

@@ 7 Conclusion
Finally, from a theoretical standpoint, we have discussed a type of communication that — we believe — is somewhat overlooked by the economic literature: the sort of persuasion that comes from the tangible results of economic actions, rather than from information design by a sender or the implicit informational content of a given equilibrium behavior. We believe that a deeper investigation of this force could help explain a variety of phenomena that go beyond productive incentives, including voter polarization, excessive debt accumulation, and other policy distortions.
@@ A.1 Examples
The characteristics of the views of example 3 — in particular, the supermodularity of expected output and noise log-concavity — recall the returns structure of Heidhues et al., (2018).
@@ A.3 Benchmark: Unawareness of Disagreement and Myopia
Next, consider part (ii). Let players be unaware of disagreement. Fix $i\in\{A,B\}$ and $m_{1}^{i}\in M$. By unawareness of disagreement, player $i$ plays as if the initial models of the two players were $(m^{i}_{1},m^{i}_{1})$. Since the two hypotheses compared by player $i$ at the end of $t=1$ coincide — the alternative model is the same as the null model — it has $m^{i}_{1}=m^{i}_{2}$ with certainty. Hence, if $m^{i}_{1}=\mathcal{H}$ then, $i$’s behavior is pinned down by 6. If $m^{i}_{1}=\mathcal{L}$ then, $i$’s behavior is pinned down by 7. It follows that
@@ A.4 Other Proofs
Consider now the case where there exists $k\in\{k^{A},k^{B}\}$ such that $m^{A}_{k}\neq m^{B}_{k}$. Without loss of generality, assume $k^{A}$ is such that $m^{A}_{k^{A}}\neq m^{B}_{k^{A}}$. Fix effort $e^{A},e^{B}\in\mathcal{E}$. $\Pi_{(e^{A},k^{A})}=(m^{A}_{k^{A}}(\cdot|e^{A}),m^{B}_{k^{A}}(\cdot|e^{A}),\mathcal{Y})$ is an informative experiment for states $m^{A}$ and $m^{B}$. There are two possible cases. First, if $m^{A}_{k^{B}}=m^{B}_{k^{B}}$, then $(y^{B},e^{B})$ are uninformative when it comes to discriminating between $m^{A}$ and $m^{B}$. Because $y^{A}$ and $y^{B}$ are independent, test LR will only use information about $(y^{A},e^{A},k^{A})$. Hence, we have $\phi_{m^{A}}^{m^{B}}(e^{A},e^{B},k^{A},k^{B}|m^{A})=\phi_{m^{A}_{k^{A}}}^{m^{B}_{k^{A}}}(e^{A},k^{A}|m^{A}_{k^{A}})$ which is increasing in $e^{A}$ by lemma A.1. The second case is $m^{A}_{k^{B}}\neq m^{B}_{k^{B}}$, in which case $\Pi_{(e^{B},k^{B})}=(m^{A}_{k^{B}}(\cdot|e^{B}),m^{B}_{k^{B}}(\cdot|e^{B}),\mathcal{Y})$ is also an informative experiment with states $m^{A}$ and $m^{B}$. Assume that this is the case and define the composite experiment $\Pi_{(e^{A},k^{A},e^{B},k^{B})}=(\Pi_{(e^{A},k^{A})},\Pi_{(e^{B},k^{B})})$. Next, consider the composite experiment $\Pi_{(\hat{e}^{A},k^{A},e^{B},k^{B})}=(\Pi_{(\hat{e}^{A},k^{A})},\Pi_{(e^{B},k^{B})})$, where $\hat{e}^{A}>e^{A}$. By assumption 3, $\Pi_{(\hat{e}^{A},k^{A})}$ is more informative than $\Pi_{(e^{A},k^{A})}$. Since $\Pi_{(\hat{e}^{A},k^{A})}$, $\Pi_{(e^{A},k^{A})}$ and $\Pi_{(e^{B},k^{B})}$ are independent, $\Pi_{(\hat{e}^{A},k^{A},e^{B},k^{B})}$ is more informative than $\Pi_{(e^{A},k^{A},e^{B},k^{B})}$ — the proof is analogous to the one provided for lemma 4.1. Hence, by lemma A.1, $\phi_{m^{A}}^{m^{B}}(\hat{e}^{A},e^{B},k^{A},k^{B}|m^{A})\geq\phi_{m^{A}}^{m^{B}}(e^{A},e^{B},k^{A},k^{B}|m^{A})$, which proves that $\phi_{m^{A}}^{m^{B}}(e^{A},e^{B},k^{A},k^{B}|m^{A})$ is increasing in $e^{A}$. The proof that $\phi_{m^{A}}^{m^{B}}(e^{A},e^{B},k^{A},k^{B}|m^{A})$ is increasing in $e^{B}$ is a replication of the same argument. The proof that $\phi_{m^{B}}^{m^{A}}(e^{A},e^{B},k^{A},k^{B}|m^{B})$ is increasing in $e^{A}$ and $e^{B}$ is analogous.
@@
Next, note that, if Ann and Bob’s initial models are $(\mathcal{H}_{x},\mathcal{L}_{y})$ and $(\mathcal{L}_{x},\mathcal{H}_{y})$, at $t=2$ both players will certainly have one view of type $\mathcal{H}$ — although they might still hold such view for different technologies. This means that, at $t=2$, both players will choose effort $e^{\mathcal{H}}$. Hence, by strict FOSD monotonicity and $Q_{x}=Q_{y}$,
@@
Next, we prove part (ii). Fix any equilibrium $s\in\hat{S}$. Let $m\in M$. By the standard arguments of the previous proposition, when players start the game agreeing on $m$, in equilibrium they will exert at most effort $e^{\mathcal{H}}$ in every period. Since models $(\mathcal{H},\mathcal{L})$ and $(\mathcal{L},\mathcal{H})$ are both correct with probability $\frac{1}{2}$, and effort and technology choices are only contingent on models — not changed between periods — it holds $\mathbb{E}_{p}[Y_{s}(m,m,Q)]\leq 2(y^{\mathcal{H}}_{\mathcal{H}}+y^{\mathcal{H}}_{\mathcal{L}})$.
Now, consider the disagreeing team with $\mathbf{m}_{1}=((\mathcal{H}_{x},\mathcal{L}_{y}),(\mathcal{L}_{x},\mathcal{H}_{y}))$. By our previous results, $s$ must be such that, for $i=A,B$, $s^{i}_{1}((\mathcal{H}_{x},\mathcal{L}_{y}),(\mathcal{L}_{x},\mathcal{H}_{y}))=e^{\mathcal{H}}+\hat{\Delta}_{e}$ for some $\hat{\Delta}_{e}\geq 0$.
@@
$Y_{s}(m^{A}_{1},m^{B}_{1},Q)=\mathbb{E}_{\mathcal{H}}[Y|e^{\mathcal{H}}+\hat{\Delta}_{e},x]+\mathbb{E}_{\mathcal{L}}[Y|e^{\mathcal{H}}+\hat{\Delta}_{e},y]+y^{\mathcal{H}}_{\mathcal{H}}$
@@
$Y_{s}(m^{A}_{1},m^{B}_{1},Q)=\mathbb{E}_{\mathcal{H}}[Y|e^{\mathcal{H}}+\hat{\Delta}_{e},y]+\mathbb{E}_{\mathcal{L}}[Y|e^{\mathcal{H}}+\hat{\Delta}_{e},x]+y^{\mathcal{H}}_{\mathcal{H}}$
@@
$\mathbb{E}_{p}[Y_{s}(m^{A}_{1},m^{B}_{1},Q)]=y^{\mathcal{H}}_{\mathcal{H}}+y^{\mathcal{H}}_{\mathcal{L}}+\mathbb{E}_{\mathcal{H}}[Y|e^{\mathcal{H}}+\hat{\Delta}_{e},x]+\mathbb{E}_{\mathcal{L}}[Y|e^{\mathcal{H}}+\hat{\Delta}_{e},y]$
@@
$\mathbb{E}_{p}[Y_{s}(m^{A}_{1},m^{B}_{1},Q)]-\mathbb{E}_{p}[Y_{s}(m,m,Q)]\geq\underbrace{(\mathbb{E}_{\mathcal{H}}[Y|e^{\mathcal{H}}+\hat{\Delta}_{e},x]-y^{\mathcal{H}}_{\mathcal{H}})}_{\geq 0}$